what volume (in mL) of 18.0M H2SO4 is needed to contain 2.45g H2SO4

1
Expert's answer
2022-05-16T17:02:03-0400

1. Using the formula n = m / M , we find the substance quantity (n) of 2.45 g h2so4. M is the molar mass of H2SO4 .

n=m/M= 2.45/98=0.025 mol

2. From the formula CM= n /Vwe derive the formula for finding V: V = n / CM

V=n/CM = 0.025/18= 0.001388889 litr

To find the milliliter, multiply the liter by 1000: 0.001388889*1000=1.38889 ml ≈ 1.4 ml

ANSWER: 1.388889 ml or approx 1.4 ml